OpenStudy (anonymous):

What am I doing wrong? solve for x: 1+4e^(x/2)=3. 4e^(x/2)=2 e^x/2=.5 x/2=ln.5 x=2ln.5 answer: -2ln2

OpenStudy (anonymous):

The short answer is nothing, the expressions are equivalent. I suppose the integer one is marginally prettier.

OpenStudy (anonymous):

1+4e^(x/2)=3 4e^(x/2)=2 e^(x/2)=1/2 lne^(x/2)=ln(1/2) x/2=ln1-ln2 x/2=0-lnw x=-2ln2

OpenStudy (anonymous):

x=2ln(1/2) = 2(ln 1 -ln 2) = -2 ln 2